OpenStudy (anonymous):

7.Create your own polynomial with a degree greater than 2 and find the zeroes of the function

OpenStudy (anonymous):

x^4-2 x^3-25 x^2+26 x+120

OpenStudy (anonymous):

start with \[(x-1)(x+1)(x+2)\] multiply out and then say the zeros are-2,-1 and 1

OpenStudy (zarkon):

\[x^3\]

OpenStudy (anonymous):

when you multiply you get \[x^3+2 x^2-x-2\]

OpenStudy (anonymous):

i like zarkon's answer more.

OpenStudy (anonymous):

\[P(x) \in \left\{\prod_{i=1}^n(a_ix - b_i) \| n \in Z_+, n > 2 \right\}\]\[Z = \left\{\frac{b_i}{a_i} \| 1 \leq i \leq n \right\} \]

OpenStudy (zarkon):

that was going to be my second choice ;)